OpenStudy (anonymous):

I need help with this - I know I need to multiple by 100 to get ride of the decimals solve by the elimination method 0.5x-0.2y=4 0.5x+0.3y=-7/17 3x-2y=40 5x+3y=-70/17 Then I am lost - can some help me?

OpenStudy (anonymous):

after multiplying with 10, the first equation should be: \[5x - 2y = 40\] probably just a typo, tho...

OpenStudy (anonymous):

you are right - typing to fast - the problem is 0.3x-0.2y=4 0.5x+0.3y=-7/17

OpenStudy (anonymous):

multiply by 100

OpenStudy (anonymous):

why by 100?

OpenStudy (anonymous):

3x-2y=40 and 5x+3y=-70/17 you do this to get rid of the decimals

OpenStudy (anonymous):

that's correct (altough unnessesary). But this was a multiplication with 10! 0.3x multiplied with 100 would be 30x !

OpenStudy (anonymous):

Now - 'elimination method' means, multiply one equation with some factor, so that, after adding both of the equations together either x or y is eliminated... Can you come up with something on your own?

OpenStudy (anonymous):

You are right I went back to notes - for elimination method it multiplication of 10 not 100